OpenStudy (anonymous):

Show that x=x1+(x2-x1)t, y=y1+(y2-y1)t, -∞

OpenStudy (anonymous):

it is clearly linear...now replace t with 0 and 1 to show that the two points are hit.

OpenStudy (anonymous):

@amistre64 Can u pls suggest how to go about this one.....

OpenStudy (amistre64):

define a direction vector between the points

OpenStudy (amistre64):

given a point (a,b), and a vector <p,q> the parametric line formed by then is simple: x = a + pt y = b + qt

OpenStudy (anonymous):

I have used the method suggested by "Math???" but am not sure if it is right. We have not been taught to use vectors in solving such questions

OpenStudy (amistre64):

or solve for t such that: t = (x-a)/p = (y-b)/q

OpenStudy (amistre64):

it would be good to know what you HAVE been taught to put this into context, otherwise there are a myriad of different ways to approach it

OpenStudy (amistre64):

another method would be to solve for t in terms of x, and sub that into the t in y to undo the parameter

OpenStudy (amistre64):

x = a + pt, t = (x-a)/p y = b + qt = b + q(x-a)/p

OpenStudy (anonymous):

We are using the book "Calculus - Graphic, Numerical and Algebraic" by Finney, Demana, Waits and Kennedy

OpenStudy (anonymous):

We have been taught to solve by either putting in values of by expressing both in terms of t and then equating them and getting a final equation

OpenStudy (amistre64):

define the point slope of the line between the given points, and compare that to the undoing of the parameter setup; if they are equal, they are the same ....

OpenStudy (anonymous):

can u pls show how this is to be done.....i am struck with only this one. I have done the rest of the homework...

OpenStudy (anonymous):

make 't' the subject in both equations and equate them. You'll get the equation of the line.

OpenStudy (anonymous):

@kelvinltr Can u pls show the actual working

OpenStudy (amistre64):

i did show the work ...

OpenStudy (amistre64):

it would be helpful if you could get your hands dirty by either attempting what has been suggested, or asking question about what might be confusing you about it.

OpenStudy (anonymous):

equation of a line passing through (x1,y1) and (x2,y2) is \[\frac{ x-x1 }{ x2-x1 }=\frac{ y-y1 }{ y2-y1 }\\\] let \[t=\frac{ x-x1 }{ x2-x1 }=\frac{ y-y1 }{ y2-y1 }\\\] then make x and y the subjects of the two equations to get the equations given in the sum

OpenStudy (anonymous):

I did convert both in terms of t and got the first equation you have written \[t=\frac{ x-x1 }{ x2-x1 } and t=\frac{ y-y1 }{ y2-y1 }\] From this I got \[\frac{ x-x1 }{ x2-x1 }=\frac{ y-y1 }{ y2-y1 }\]

OpenStudy (anonymous):

Now what to do further???

OpenStudy (amistre64):

the slope from (x1,y1) to (x2,y2) can defined to be (y2-y1)/(x2-x1) point slope form of a line is then: (y-a) = (y2-y1)/(x2-x1) (x-b) ; for any point (a,b); choose the point (x1,y1) (y-y1) = (y2-y1)/(x2-x1) * (x-x2) equate that to the undoing of the parameters: t = (x-x1)/(x2-x1) y = y1 + (y2-y1)/(x2-x1) * (x-x1) (y-y1) = (y2-y1)/(x2-x1) * (x-x1)

OpenStudy (amistre64):

might have a typo in there :)

OpenStudy (anonymous):

@amistre64 @kelvinltr thanks for your help.!! Finally by adding up all ideas and working, I think I got the answer in the form I wanted. Now only my teacher will tell if it is right or not....Thanks again!!

OpenStudy (amistre64):

:) good luck